OpenStudy (anonymous):

Solve for g. √6g-23=√12-g

OpenStudy (anonymous):

how will i solve g

OpenStudy (callisto):

Is your question like this? \[\sqrt{6} g -23 = \sqrt{12}-g\]

OpenStudy (experimentx):

In this is your equation than no solutions exist from wolfam http://www.wolframalpha.com/input/?i=sqrt%286g%29+-+23+%3D+sqrt%2812-g%29

OpenStudy (callisto):

@samifarooqi What is your question exactly? 1. \[\sqrt{6} g -23 = \sqrt{12} -g\]2. \[\sqrt{6g -23} = \sqrt{12 -g}\] Or?

OpenStudy (anonymous):

liek that

OpenStudy (callisto):

Okay, \[\sqrt{6g-23} = \sqrt{12-g}\] square both sides \[\sqrt{6g-23}^2 = \sqrt{12-g}^2\] Then you'll get \[6g-23 = 12-g\] Do you know how to do it from here?

OpenStudy (anonymous):

no....i have trouble in math

OpenStudy (callisto):

But do you understand the above workings?

OpenStudy (anonymous):

o yes i got it thanks

OpenStudy (callisto):

You can solve it?